# Public Infrastructure Investments for Space Market Development

### v1

@@ 5 Conclusion
This essay addresses a fundamental challenge facing public space agencies: how to allocate limited budgets between direct purchases, direct transfers, and shared infrastructure investments to develop and sustain markets for advanced technology systems. The core economic insight—that shared infrastructure creates non-rival benefits that can sustain more competitors per marginal dollar than direct support mechanisms—follows from standard principles of public economics (Samuelson, 1954). The commercial space station application demonstrates this: under plausible conditions, public investment in shared core infrastructure enables two competing firms to earn substantial profits, while allocating that budget entirely toward direct purchases alone results in industry-wide losses. The sustainable competition diagram makes these kinds of effects transparent to decision-makers.
This framework applies broadly to advanced technology domains with high fixed costs, uncertain market demand, non-market motivations, and separable system functions enabling shared infrastructure. Public science and technology agencies in such sectors face similar tensions between procurement objectives and market development goals, while private investors must evaluate opportunities where government decisions can matter more than market fundamentals. Recent experience provides sobering evidence of these constraints: Collins Aerospace’s withdrawal from NASA’s xEVAS contract shows how market forces can overwhelm program design choices.
As global investment in advanced technologies accelerates, frameworks that help planners systematically compare economic architectures will become more valuable to organizations seeking to develop markets for these systems. However, significant shares of demand for these systems are often driven by non-economic motivations. Analyses focused on first-best regulatory pathways to maximize economic welfare rather than second-best investment strategies to support particular goals can miss opportunities for policy relevance. Market development plans require theoretical and empirical foundations that make economic analysis credible and actionable for the organizations that must implement them.

### v2

@@ 5 Conclusion
This paper addresses a fundamental challenge facing public space agencies: how to allocate limited budgets between direct purchases, direct transfers, and shared infrastructure investments to develop and sustain markets for advanced technology systems. The core economic insight—that shared infrastructure creates non-rival benefits that can sustain more competitors than direct support mechanisms—follows from standard principles of public economics (Samuelson, 1954). The crewed space station application demonstrates this: under plausible conditions, public investment in a shared core module enables two competing firms to earn substantial profits, while allocating that budget entirely toward direct purchases results in industry-wide losses. The sustainable competition diagram makes these trade-offs transparent to decision-makers.
The diagram can be applied broadly to advanced technology domains with high fixed costs, uncertain external demand, non-economic motivations, and separable system functions enabling shared infrastructure. Public science and technology agencies in such domains may face similar tensions between procurement objectives and industrial base or market development goals, while private investors must evaluate opportunities where government decisions can matter more than market fundamentals. On the other hand, the scenario analyses reveal that economic conditions can dominate program design choices in determining market outcomes. This highlights both the potential and the limits of public infrastructure investments for market development.
This logic can be extended to international cooperation. Infrastructure shared across partners reduces duplicated fixed costs, implying efficiency gains from cooperation. The Lunar Gateway exemplifies this: NASA provides power and propulsion, ESA contributes habitation and refueling modules, JAXA provides life support systems, Canada contributes a robotic arm, and the UAE provides the crew airlock, distributing development costs while enabling all partners to access common capabilities (NASA, 2024a). Yet national interest in space systems is often driven by non-economic motivations, and geopolitical considerations can dominate efficiency calculations. China developed Tiangong independently after the Wolf Amendment effectively excluded Chinese participation in the ISS. Europe developed an independent satellite navigation system, Galileo, despite GPS being freely available. Fragmentation of infrastructure investment can also fragment external demand, since partners in shared infrastructure can become natural customers for each other’s commercial providers. While economic analysis may not be able to resolve whether efficiency gains from cooperation outweigh the non-economic benefits planners perceive from strategic autonomy, it can clarify the economic stakes of those choices.
Market development plans require economic analysis grounded in the decisions planners actually face. Crewed space stations are a convenient example but not the only one. As global investment in advanced technologies accelerates, frameworks that help planners assess economic trade-offs and uncertainty will become more valuable to public and private technology investors.